Jim is trying to determine the height of a flagpole using shadows. Jim is 6 feet tall and his shadow is 9 feet long. At the same time of day, and at the same position, the flagpole's shadow is 60 feet long. How tall is the flagpole?
Answer by checkley79(3341)

6/8=x/60 cross multiply
8x=6*60
8x=360
x=360/8
x=45 feet tall for the flagpole.
Proof:
6/8=45/60
3/4=3/4